OpenStudy (anonymous):

Find the limit of x^x^2 as x approaches to 0.

zepdrix (zepdrix):

\[\Large \lim_{x\to0}x^{(x^2)}\] Let's use this familiar identity to rewrite our expression,\[\Large x^{(x^2)} \qquad=\qquad e^{\ln[x^{(x^2)}]}\]We can pass the limit into the exponential,\[\Large e^{\lim_{x\to0}\ln[x^{(x^2)}]}\]Ignoring the exponential base, let's solve the limit,\[\Large \lim_{x\to0}\ln[x^{(x^2)}] \qquad=\qquad \lim_{x\to0}(x^2)\ln x\] From here we're getting the indeterminate form \[\Large 0\cdot-\infty\]. We want to get it into a form where we can apply L'Hopital's Rule. This might seem a little tricky but we can rewrite our x^2 as, \[\Large x^2=\frac{1}{\left(\frac{1}{x^2}\right)}\] Which makes our limit become, \[\Large lim_{x\to0}\frac{\ln x}{\frac{1}{x^2}}\]Now our limit is approaching the indeterminate form, \[\Large \frac{-\infty}{\infty}\] Yay! We can apply L'Hopital's! :)

zepdrix (zepdrix):

Applying L'Hop gives us, \[\Large lim_{x\to0} \frac{\left(\dfrac{1}{x}\right)}{\left(-2\dfrac{1}{x^3}\right)}\]Which we can write as, \[\Large lim_{x\to0} \frac{\left(\dfrac{1}{x}\right)}{\left(-2\dfrac{1}{x^2}\right)\dfrac{1}{x}}\] And then cancel some stuff, \[\Large lim_{x\to0}\frac{1}{-2\dfrac{1}{x^2}}\] Which we can rewrite as, \[\Large lim_{x\to0}-\frac{1}{2}x^2\]

zepdrix (zepdrix):

So now we can see that as x approaches 0, our function is also approaching zero. BUT REMEMBER! This was all inside of our exponent! \[\Large \lim_{x\to0}-\frac{1}{2}x^2=0\]

zepdrix (zepdrix):

\[\Huge e^{\lim_{x\to0}\ln[x^{(x^2)}]} \qquad=\qquad e^0\]